OpenStudy (anonymous):

The circumference of a circle that has a central angle of 50o that intercepts an arc of length 15 cm is _____ cm.

OpenStudy (nurali):

Solve 50/360=15/c for c. The solution is c=360*15/50=100 cm

OpenStudy (anonymous):

@nurali can you help me with my question